### v1

# Meta-generalized-gradient exchange-correlation functionals for spin-orbit coupled electrons

Abstract
The prominence of density functional theory (DFT) in the field of electronic structure computation stems from its ability to usefully balance accuracy and computational effort. At the base of this ability is a functional of the electron density: the exchange-correlation (xc) energy. This functional satisfies known exact conditions that guide the derivation of approximations. The strongly-constrained-appropriately-normed (SCAN) approximation stands out as a successful, modern, example. In the presence of spin-orbit coupling, however, the xc energy must depend not only on the electron density but also on spin currents – a fact which is not captured by SCAN or by any mainstream functional approximation. In this work, we demonstrate how the SU(2) gauge-invariance of the exact functional allows us to add the necessary dependence on spin currents in the SCAN functional (here called JSCAN) — and similar meta-generalized-gradient functional approximations — in a non-empirical manner. The prescription further ensures that the conjugate xc-fields of the spin currents have vanishing local torques as in the exact case. In passing, a spin-current dependent generalization of the electron localization function (here called JELF) is also derived. The extended forms are implemented in a developer’s version of the Crystal23 program. Applications on molecules and materials confirm the practical relevance of the extensions.
pacs:
71.15.Mb, 71.15Rf, 31.15.E-Introduction. Relativistic effects play an important role in our understanding of the electronic structure of matter. Greiner (2012) Spin-orbit coupling (SOC), in particular, emerges naturally in the two-component Pauli equation (a simplified low-energy version of the Dirac equation) and plays a crucial role in our understanding of the band structure of materials such as Si, Ge and GaAs, which powered the semiconductor revolution. Yu and Cardona (2010) More recently, the interest in spin-orbit interactions in condensed matter has grown even further following the discovery of a host of materials which exhibit nontrivial topological properties Bernevig and Hughes (2013); Vanderbilt (2018) and the explosion of interest in new techniques to accurately control the spin of the electrons in spintronic and quantum computing devices.utić et al. (2004)
Just as an electric potential couples to the particle density and a magnetic field couples to the spin and to the particle (i.e. ordinary) current, SOC can be described as a tensor field which couples to the spin current (SC), i.e., the tensor product of spin and ordinary current. Unlike the ordinary current, the spin current is invariant under time reversal (being the product of two time-reversal odd quantities) and therefore can freely flow in the time-reversal invariant ground state of a non magnetic system.Tokatly (2008); Droghetti et al. (2022)
It is sometimes adequate (for systems composed of light atoms) to include SOC as a small perturbation on top of the Kohn-Sham (KS) potential of ordinary density functional theory (DFT).Kohn (1999) This approach can in principle find the particle density and total energy of real interacting ground states by solving a single-particle problem in the presence of a self-consistent exchange-correlation (xc) potential, which is a functional of the density. Also, the eigenvalues of the KS single-particle equation provide a useful representation of the band structure and related properties.Seidl et al. (1996); Perdew et al. (2017); Nazarov and Vignale (2011); Prokopiou et al. (2022)
However, there is no reason why a density functional approximation that was developed for systems lacking SOC would still yield accurate results when strong SOC are present. To accomplish this more ambitious task, we need an exchange-correlation energy functional that depends not only on the density but also — and crucially — on the spin current density, and does so in a universal (i.e., system-independent) manner, as mandated by the original spin-current density functional theory (SCDFT)Vignale and Rasolt (1988); Bencheikh (2003) and its more recent developments. Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2023a) In practice, this means including in the effective potential, not only the “bare” SOC, but also its “dressing” due to many-body effects, in analogy to the dressing of the bare electron-nuclear potential by Hartree and exchange-correlation (xc) effects.
Crucially, any xc energy functional that depends on the spin current density must be invariant — not only under U(1) gauge transformationsVignale and Rasolt (1987); Tao and Perdew (2005) — but also under SU(2) transformationsVignale and Rasolt (1988); Bencheikh (2003); Pittalis et al. (2017). The latter can be viewed as local rotations of the spinorial state of the electronsFröhlich and Studer (1993); Jin et al. (2006). Failing to satisfy SU(2) invariance would have disastrous consequences for a density functional approximation: it would mean that the approximation does not distinguish between actual variations of the spin current density — which can have physical consequences — and variations arising from gauge transformations, which must have no consequence! Thus the gauge invariance of the exact xc-energy functional acts as a most important exact constraint, which guides the construction of approximate functionals.
While different approximate functionals obviously have different strengths and weaknesses, they all must satisfy the constraint of gauge-invariance. Hence, we focus on the strongly-constrained-appropriately-normed (SCAN) functional,Sun et al. (2015) which is currently recognized as the least empirical general-purpose approximation in the class of Meta-Generalized-Gradient Approximations (MGGA).Perdew et al. (2009); Sun et al. (2013, 2015, 2016); Tran et al. (2016); Peng et al. (2016); Brandenburg et al. (2016); Isaacs and Wolverton (2018); Tran et al. (2019) However, our procedure is applicable to any approximate functional in this class.
The exceptional performance of the SCAN is fundamentally due to its ability to satisfy altogether a large number of known exact conditions for the xc-energy functional of DFT. Furthermore, among its crucial ingredients, the KS kinetic-energy density appears in a combination of variables that was originally proposed for the so-called electron localization function (ELF) – a quantity that provides a vivid visualization of molecular bonds and atomic shells (more below).Becke and Edgecombe (1990a); Silvi and Savin (1994)
But DFT (and thus SCAN) cannot deal with non-perturbative SOC! This brings forth the necessity to switch to SCDFT both in principleVignale and Rasolt (1988); Bencheikh (2003) and in practice.Rohra and Görling (2006); Heaton-Burgess et al. (2007); Trushin and Görling (2018); Desmarais et al. (2023a, 2020a); Bodo et al. (2022); Comaskey et al. (2022) The dependence of the xc energy functional on spin currents in SCDFT is strongly constrained by the requirement that its form remains invariant under local SU(2) gauge transformations (defined below). Taking as a starting point the SCAN and the ELF forms, in this work we show how to make them functionals of the spin currents while preserving the fundamental SU(2) gauge invariance of the theory. The resulting generalizations are here called JELF and JSCAN. A set of computed results for molecules and materials demonstrates the practical usefulness of our non-empirical extension of the SCAN functional. An implementation is presented in a developer’s version of the Crystal23 program.Erba et al. (2023)
SCDFT in a nutshell. We shall consider ground states at geometries that do not magnetize nor host particle currents spontaneously. Furthermore, magnetic couplings (Zeeman field, Abelian vector potential) that break time-reversal symmetry are not included. Accounting for SOC, however, we must consider together with the particle density, , the spin currents, , and corresponding external couplings: a scalar, , and a non-Abelian vector potential, . The former accounts for the electron-nuclear potential, while the latter provides a useful formal representation of SOC.Fröhlich and Studer (1993); Bencheikh (2003); Abedinpour et al. (2010); Zhang and Vignale (2016); Droghetti et al. (2022) The energy density functional of interest, thus, reads as follows:
where is the kinetic energy of the KS system, while and are Hartree and xc energies. Here and in the following, we denote with bold characters, , quantities with spatial indices (Greek lower indices, , when written explicitly); and use an arrow, , to denote quantities with spin indices (upper Latin indices, , when written explicitly) — and have values ,,. Therefore and have both spatial and spin indices. Contractions over spatial indices are denoted with “” and Einstein convention is used for summing over repeated indices. Unless otherwise stated, we use Hartree atomic units in which . The corresponding KS equations are:
where , with being the vector of Pauli matrices, is a Hartree-xc-scalar potential (analogous to the Hxc-scalar potential of DFT) and is a non-Abelian xc-vector potential (which is specific to SCDFT). The occupied (lowest in energy) two-component KS spinors allow us to compute the particle density, , and the spin current of the real state and, thus, the energy as well. All this is exact in principle, while, in practice, we are bound to approximate .
Invoking a MGGA, it is convenient to consider the so-called generalized-KS approach to SCDFT.Desmarais et al. (2023a) This means that for generating the single-particle equations, we differentiate the energy functional expression explicitly w.r.t. the orbitals rather than w.r.t. to the densities not (a) — nowadays, this has become a standard procedure in DFT.Neumann et al. (1996)
A central property of the xc-energy functional in SCDFT is its form invariance under local SU(2) transformations, i.e., transformations that act on the KS spinors as follows: , where , i.e., subjecting them to different rotations at different points in space. This invariance is expressed by the equation
where , , is a matrix describing a rotation in around the direction by an angle , and . not (b)
It is important to notice that the spin current itself is not a gauge invariant quantity since it can appear as a consequence of a generic SU(2) transformation acting on a state in which . The invariance of stated in Eq. (3), implies that only certain gauge-invariant combinations of with other KS quantities can legitimately appear in . We now tackle the task of generalizing the SCAN functional (and the closely related ELF) to include a dependence on spin currents that complies with Eq. (3).
Extension of SCAN and ELF to spin-current carrying states. The SCAN functional setsSun et al. (2015) where interpolates between the semi-local energy densities for single-orbital densities () and for slowly-varying densities (), which only depend on (a gauge-invariant quantity). The interpolation function is controlled by the variable . The latter quantity entails three kinetic energy densities: the positive-definite conventional defined in terms of the occupied (one-component) KS orbitals {}, the von Weizsäcker — that is the bosonic expression of — where , and the kinetic energy of the noninteracting Fermi gas at uniform density : .
The success of the SCAN is explained not only by the fact that it satisfies a large set of known exact conditions but also by the fact that it uses, specifically, as a key variable. The tends to zero in the iso-orbital limit (i.e., when the density is dominated by a single occupied orbital), tends to one in the uniform-density limit, and tends to infinity in regions dominated by density overlap between closed shells. In fact, is a useful building block of density functional approximations,Becke (1996a); Perdew et al. (2009); Sun et al. (2013) and it is the main ingredient of the electron localization function,Becke and Edgecombe (1990b); Silvi and Savin (1994); Savin et al. (1997) – a simple but effective descriptor of molecular bonds and atomic shells.
More in general, the kinetic energy density must be expressed in terms of the occupied two-component KS spinors . Crucially, let us next consider the transformation of the kinetic energy under a local SU(2) transformation, , of the spinors. For the states considered in this work, one readily finds: Pittalis et al. (2017) ; for the definition of see below Eq. (3). As a consequence of which, it is apparent that will change, too. Hence, neither the SCAN functional nor the ELF are SU(2)-form invariant. Such an invariance, however, is not only an exact mathematical property of the xc-energy functional in SCDFT but — because a local SU(2) transformation may be regarded as part of an overall gauge transformation (i.e., a change in the description which cannot change the physics) — it is also a proper feature for any quantity that should carry direct chemical-physical information. Fortunately, in one stroke, the substitution , solves both issues: i.e., it introduces an explicit dependence on the spin currents and enforces the sought form invariance. In detail,
implies
and
Note, the above substitution should not be confused with the analogous well-known substitution that involves the (paramagnetic) particle current:Tao and Perdew (2005) The latter derives from the consideration of local U(1) transformations. The importance of the dependence of density functionals on the particle current has been largely demonstrated. Dobson (1993); Becke (1996b, 2002); Maximoff et al. (2004); Burnus et al. (2005); Tao and Perdew (2005); Pittalis et al. (2007, 2009); Räsänen et al. (2009); Oliveira et al. (2010); Zhu et al. (2016); James W. Furness and Teale (2016); Holzer et al. (2022)
Next, from Eq. (5), taking a functional derivative with respect to , we obtain
where denotes the derivative of wrt its own argument. This is a non-Abelian effective connection; i.e., an xc-spin-vector potential. This expression shows that is parallel to at each point in space and, therefore,
For the non-magnetic ground states considered in this work, the JSCAN satisfies the non-Abelian torque conditions exactly. not (c)
In summary, Eqs. (4)-(8) are the key equations of this work — Eqs. (4), (7) and (8) apply to any MGGA which uses through (for example the TASK functional).Aschebrock and Kümmel (2019)
Applications. We have implemented a self-consistent treatment of the JSCAN functional, including effective non-Abelian vector potentials, as well as the JELF in a developer’s version of the Crystal23 code.Erba et al. (2023) Computational and implementation details are provided in the supporting informationESI (see also Refs. Desmarais et al., 2021a, 2020b, 2019a, 2019b; Metz et al., 2000; Peterson et al., 2003, 2007; Doll et al., 2001; Doll, 2001; Doll et al., 2006; Civalleri et al., 2001; Laun et al., 2018; Heyd et al., 2005; Lebedev, 1976, 1977; Towler et al., 1996; Lehtola et al., 2020 therein). In the calculations, we employ the restored-regularized r-SCAN, rather than the original SCAN functional, to improve numerical stability and performance.Furness et al. (2022)
Exx-Only | Exx & Corr | |||
J-rSCAN | HF | J-rSCAN | RPAx | |
FI | 0.248 (0.225) | 0.322 | 0.232 (0.222) | 0.360 |
HI | 0.337 (0.276) | 0.343 | 0.324 (0.281) | 0.343 |
I | 0.333 (0.276) | 0.364 | 0.317 (0.278) | 0.396 |
HAt | 0.927 (0.675) | 0.891 | 0.884 (0.213) | 0.887 |
We apply our SU(2) gauge-invariant J-rSCAN approach to molecules on reproducing the SOC-induced modification of the valence electronic structure in halogen diatoms and hydrides, which have served as model systems in previous studies.Desmarais et al. (2021b, c, 2023b) Values of the SOC-induced modification to the HOMO-LUMO gap are reported in Table 1, first at the exchange-only level of theory, and then including effects of exchange and correlation. Values are presented with exchange-only (gauge-invariant) J-rSCAN functional, and numbers in parentheses represent corresponding results with the standard (gauge-dependent) rSCAN approach.
Inclusion of the spin currents in the functional greatly improves the agreement with Hartree-Fock (HF) results. In the case of the HI molecule, for example, the HOMO-LUMO gap is increased from 0.276 eV to 0.337 eV; which reduces the relative error from 20% to 2%. In the case of HAt, the effect of SOC on the HOMO-LUMO gap is increased from 0.675 eV to 0.927 eV; relative error decreased from 24% to 4%.
At the xc level of theory, values are benchmarked against the random phase approximation including the dynamical response of Fock exchange (RPAx).Ullrich (2011) In detail, this comparison assumes that the effect of SOC on the fundamental gap is to a good approximation the same as the effect of SOC on the optical gap. In all cases, inclusion of currents in the xc model improves agreement against the RPAx. The most extreme case is HAt where the effect of SOC on the gap is increased from 0.213 eV to 0.884 eV by inclusion of spin-currents in the MGGA, essentially coinciding with the RPAx value of 0.887 eV.
Next, we discuss electron localization through the JELF [see Eq. (6)]. Let us begin by recalling that large values of , corresponding to a (J)ELF close to one, indicate that electrons are highly localized. Here we look at the effect of SOC on the (J)ELF. For the I molecule, after a self-consistent solution with J-rSCAN, including both exchange and correlation effects, in Fig. 1(a) we plot the ELF and JELF indicators as differences w.r.t. the calculation without SOC. Yellow contours indicate that SOC localizes the electron. Blue contours indicate that the electron is delocalized by SOC. The leftmost panel, containing the JELF shows that electrons are localized along the bond axis by SOC. On the other hand, we observe two blue “onion rings” around the atomic centers. Thus, SOC localizes electrons along the bond and delocalizes them in the orthogonal directions. Although this analysis is consistent with previous reports,Desmarais et al. (2023c); Gulans and Draxl (2022) the degree of delocalization is largely overestimated by the gauge-dependent ELF, compared to the SU(2) form-invariant JELF, see their difference in the rightmost panel of Fig. 1(a). The JELF displays more localization on the bonding compared to the (gauge-dependent) ELF.
We proceed to apply our approach to the description of the electronic band structure of infinite periodic systems. In Table 2, we provide values of the splitting of the valence band at the high-symmetry point in molybdenum dichalcogenide Rashba systems: the inversion asymmetric hexagonal single layer MoSe and the inversion-symmetric hexagonal bulk -MoTe. The valence band structure is provided in panel (b) of Figure 1.
For both systems, about one third of the total effect of SOC on the splitting is accounted for through many-body effects by inclusion of spin currents in the functional. Low-temperature experimental values on the transport gap are available from Refs. Choi et al., 2017; Island et al., 2016; Zelewski and Kudrawiec, 2017; Conan et al., 1979, while band splittings are provided in Refs. Oliva et al., 2020; Zhang et al., 2014; Shim et al., 2014; Ross et al., 2013; Beal et al., 1972; Ruppert et al., 2014. Including spin currents in the xc form has a very small effect on the band gaps (see Table 2). On the other hand, the values of valence band splittings are considerably enhanced (from 0.14 eV to 0.20 eV in MoSe, and from 0.28 eV to 0.32 eV in MoTe), leading to an improved agreement with experiment (relative error decreased from 29% to 10% in MoSe and leading to an exact coincidence with the mean experimental value in MoTe). Overall, the accuracy of the J-rSCAN approach is comparable to more involved hybrid-GGA SCDFT treatments (which provide splittings of 0.17 and 0.32 eV, respectively for MoSe and MoTe using optimal fractions of Fock exchange) — without requiring the determination of an optimal fraction of Fock exchange:Desmarais et al. (2023a) a step that involves either empiricism or extra/external ab initio calculations.Skone et al. (2014)
2D MoSe | 3D -MoTe | |||
split | gap | split | gap | |
rSCAN | 0.14 | 1.57 | 0.28 | 0.82 |
J-rSCAN | 0.20 | 1.53 | 0.32 | 0.81 |
Exp. | 0.18 | 1.6-2.3 | 0.3-0.34 | 1.03 |
Finally, we report on the application of our SU(2) form-invariant approach to topological materials. We consider the case of Weyl semi-metals. We report on the splitting of Weyl nodes in the orthorhombic TaAs phase. We study the W1 node pair, which lies along the mirror plane. The pair of nodes is located at (in units of ), in excellent agreement with the experimental value of .Lv et al. (2015); Huang et al. (2015) Spin-current dependent terms in the xc functional account for of the total splitting along . See Fig. S1 of the supplementary material.
Conclusions. Focusing on the prominent example of the SCAN functional, we have shown how to include a dependence on spin currents in a meta-GGA xc functional form for evaluating SOC effects in such a way as to fulfill the prerequisite of SU(2) gauge invariance of the xc energy. For the states considered in this work, the current-dependent form of the approximation is furthermore strongly suggested by previous studies on the structure of the exact exchange hole (see Ref. Pittalis et al., 2017), which incidentally, also motivate the inclusion of the kinetic energy density as a variable on which the functional must depend. A large body of works have highlighted the usefulness of the kinetic energy in combination with other quantities as in the ELF (and, thus, as in the iso-orbital indicator of SCAN) as a practical means for capturing relevant local features of the many-electron state in correlated regimes beyond exchange-only effects. Of course, the deeper question of including the “right physics” beyond JSCAN remains open. Aside from exact constraints provided by the SU(2) gauge principle, the success of the benchmark calculations reported in our paper leads us to believe that we are proceeding on the right path.
Looking ahead, the approach proposed in this work will help disentangle the role of the SU(2) invariance in density functional approximations from questions more directly related to the physics of the electron-electron interaction. It is particularly appealing to explore the consequence of a full (compound) U(1)SU(2) gauge-invariance on general states that may also break time-reversal symmetry. This may help solve standing difficulties of present DFT methodologies for magnetized systems.Isaacs and Wolverton (2018); Jana et al. (2018); Ekholm et al. (2018); Fu and Singh (2018, 2019); Mejía-Rodríguez and Trickey (2019); Romero and Verstraete (2018); Buchelnikov et al. (2019); Tran et al. (2020)
Acknowledgements.
This research has received funding from the Project CH4.0 under the MUR program “Dipartimenti di Eccellenza 2023-2027” (CUP: D13C22003520001). GV was supported by the Ministry of Education, Singapore, under its Research Centre of Excellence award to the Institute for Functional Intelligent Materials (I-FIM, project No. EDUNC-33-18-279-V12).References
- Greiner (2012) W. Greiner, Relativistic Quantum Mechanics: Wave Equations (Springer Berlin Heidelberg, 2012), ISBN 9783642880827.
- Yu and Cardona (2010) P. Yu and M. Cardona, Fundamentals of Semiconductors: Physics and Materials Properties, Graduate Texts in Physics (Springer Berlin Heidelberg, 2010), ISBN 9783642007101.
- Bernevig and Hughes (2013) B. Bernevig and T. Hughes, Topological Insulators and Topological Superconductors (Princeton University Press, 2013), ISBN 9780691151755.
- Vanderbilt (2018) D. Vanderbilt, Berry Phases in Electronic Structure Theory: Electric Polarization, Orbital Magnetization and Topological Insulators (Cambridge University Press, 2018), ISBN 9781107157651.
- utić et al. (2004) I. utić, J. Fabian, and S. Das Sarma, Rev. Mod. Phys. 76, 323 (2004).
- Tokatly (2008) I. V. Tokatly, Phys. Rev. Lett. 101, 106601 (2008), URL https://link.aps.org/doi/10.1103/PhysRevLett.101.106601.
- Droghetti et al. (2022) A. Droghetti, I. Rungger, A. Rubio, and I. V. Tokatly, Phys. Rev. B 105, 024409 (2022), URL https://link.aps.org/doi/10.1103/PhysRevB.105.024409.
- Kohn (1999) W. Kohn, Rev. Mod. Phys. 71, 1253 (1999), URL https://link.aps.org/doi/10.1103/RevModPhys.71.1253.
- Seidl et al. (1996) A. Seidl, A. Görling, P. Vogl, J. A. Majewski, and M. Levy, Phys. Rev. B 53, 3764 (1996).
- Perdew et al. (2017) J. P. Perdew, W. Yang, K. Burke, Z. Yang, E. K. Gross, M. Scheffler, G. E. Scuseria, T. M. Henderson, I. Y. Zhang, A. Ruzsinszky, et al., PNAS 114, 2801 (2017).
- Nazarov and Vignale (2011) V. U. Nazarov and G. Vignale, Phys. Rev. Lett. 107, 216402 (2011), URL https://link.aps.org/doi/10.1103/PhysRevLett.107.216402.
- Prokopiou et al. (2022) G. Prokopiou, M. Hartstein, N. Govind, and L. Kronik, J. Chem. Theory Comput. 18, 2331 (2022).
- Vignale and Rasolt (1988) G. Vignale and M. Rasolt, Phys. Rev. B 37, 10685 (1988).
- Bencheikh (2003) K. Bencheikh, J. Phys. A 36, 11929 (2003).
- Rohra and Görling (2006) S. Rohra and A. Görling, Phys. Rev. Lett. 97, 013005 (2006).
- Heaton-Burgess et al. (2007) T. Heaton-Burgess, P. Ayers, and W. Yang, Phys. Rev. Lett. 98, 036403 (2007).
- Trushin and Görling (2018) E. Trushin and A. Görling, Phys. Rev. B 98, 205137 (2018).
- Desmarais et al. (2023a) J. K. Desmarais, G. Ambrogio, G. Vignale, A. Erba, and S. Pittalis, Generalized kohn-sham approach for the electronic band structure of spin-orbit coupled materials (2023a), eprint 2309.11158.
- Vignale and Rasolt (1987) G. Vignale and M. Rasolt, Phys. Rev. Lett. 59, 2360 (1987), URL http://link.aps.org/doi/10.1103/PhysRevLett.59.2360.
- Tao and Perdew (2005) J. Tao and J. P. Perdew, Phys. Rev. Lett. 95, 196403 (2005).
- Pittalis et al. (2017) S. Pittalis, G. Vignale, and F. G. Eich, Phys. Rev. B 96, 035141 (2017).
- Fröhlich and Studer (1993) J. Fröhlich and U. M. Studer, Rev. Mod. Phys. 65, 733 (1993).
- Jin et al. (2006) P.-Q. Jin, Y.-Q. Li, and F.-C. Zhang, Journal of Physics A: Mathematical and General 39, 7115 (2006).
- Sun et al. (2015) J. Sun, A. Ruzsinszky, and J. P. Perdew, Phys. Rev. Lett. 115, 036402 (2015), URL https://link.aps.org/doi/10.1103/PhysRevLett.115.036402.
- Perdew et al. (2009) J. P. Perdew, A. Ruzsinszky, G. I. Csonka, L. A. Constantin, and J. Sun, Phys. Rev. Lett. 103, 026403 (2009), URL https://link.aps.org/doi/10.1103/PhysRevLett.103.026403.
- Sun et al. (2013) J. Sun, B. Xiao, Y. Fang, R. Haunschild, P. Hao, A. Ruzsinszky, G. I. Csonka, G. E. Scuseria, and J. P. Perdew, Phys. Rev. Lett. 111, 106401 (2013), URL https://link.aps.org/doi/10.1103/PhysRevLett.111.106401.
- Sun et al. (2016) J. Sun, R. C. Remsing, Y. Zhang, Z. Sun, A. Ruzsinszky, H. Peng, Z. Yang, A. Paul, U. Waghmare, X. Wu, et al., Nat Chem 8, 831 (2016).
- Tran et al. (2016) F. Tran, J. Stelzl, and P. Blaha, The Journal of Chemical Physics 144, 204120 (2016), ISSN 0021-9606, eprint https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/1.4948636/15511401/204120_1_online.pdf, URL https://doi.org/10.1063/1.4948636.
- Peng et al. (2016) H. Peng, Z.-H. Yang, J. P. Perdew, and J. Sun, Phys. Rev. X 6, 041005 (2016), URL https://link.aps.org/doi/10.1103/PhysRevX.6.041005.
- Brandenburg et al. (2016) J. G. Brandenburg, J. E. Bates, J. Sun, and J. P. Perdew, Phys. Rev. B 94, 115144 (2016), URL https://link.aps.org/doi/10.1103/PhysRevB.94.115144.
- Isaacs and Wolverton (2018) E. B. Isaacs and C. Wolverton, Phys. Rev. Mater. 2, 063801 (2018), URL https://link.aps.org/doi/10.1103/PhysRevMaterials.2.063801.
- Tran et al. (2019) F. Tran, L. Kalantari, B. Traoré, X. Rocquefelte, and P. Blaha, Phys. Rev. Mater. 3, 063602 (2019), URL https://link.aps.org/doi/10.1103/PhysRevMaterials.3.063602.
- Becke and Edgecombe (1990a) A. D. Becke and K. E. Edgecombe, The Journal of Chemical Physics 92, 5397 (1990a), eprint http://dx.doi.org/10.1063/1.458517, URL http://dx.doi.org/10.1063/1.458517.
- Silvi and Savin (1994) B. Silvi and A. Savin, Nature 371, 683 (1994), URL https://www.nature.com/articles/371683a0.
- Desmarais et al. (2020a) J. K. Desmarais, J.-P. Flament, and A. Erba, Phys. Rev. B 102, 235118 (2020a).
- Bodo et al. (2022) F. Bodo, J. K. Desmarais, and A. Erba, Phys. Rev. B 105, 125108 (2022).
- Comaskey et al. (2022) W. P. Comaskey, F. Bodo, A. Erba, J. L. Mendoza-Cortes, and J. K. Desmarais, Phys. Rev. B 106, L201109 (2022).
- Erba et al. (2023) A. Erba, J. K. Desmarais, S. Casassa, B. Civalleri, L. Doná, I. J. Bush, B. Searle, L. Maschio, L. Edith-Daga, A. Cossard, et al., J. Chem. Theor. Comput. 19, 6891 (2023).
- Abedinpour et al. (2010) S. H. Abedinpour, G. Vignale, and I. V. Tokatly, Phys. Rev. B 81, 125123 (2010).
- Zhang and Vignale (2016) S. S.-L. Zhang and G. Vignale, Phys. Rev. B 94, 140411 (2016).
- not (a) At the level of GGAs, this implies no substantial difference because the approximate functional, however, depends on the particle density (and its gradients) explicitly. At the level of a MGGA, this implies that an extra semi-local potential shows up because of the explicit dependence on the the KS kinetic energy. This potential is usually grouped with the scalar potential Lehtola et al. (2020).
- Neumann et al. (1996) R. Neumann, R. H. Nobes, and N. C. Handy, Mol. Phys. 87, 1 (1996).
- not (b) is pure gauge vector potential, fully originated by the SU(2) transformation.
- Becke (1996a) A. D. Becke, The Journal of Chemical Physics 104, 1040 (1996a), ISSN 0021-9606, eprint https://pubs.aip.org/aip/jcp/article-pdf/104/3/1040/10779816/1040_1_online.pdf, URL https://doi.org/10.1063/1.470829.
- Becke and Edgecombe (1990b) A. D. Becke and K. E. Edgecombe, The Journal of Chemical Physics 92, 5397 (1990b), ISSN 0021-9606, eprint https://pubs.aip.org/aip/jcp/article-pdf/92/9/5397/9731640/5397_1_online.pdf, URL https://doi.org/10.1063/1.458517.
- Savin et al. (1997) A. Savin, R. Nesper, S. Wengert, and T. F. Fässler, Angewandte Chemie International Edition in English 36, 1808 (1997).
- Dobson (1993) J. F. Dobson, The Journal of Chemical Physics 98, 8870 (1993).
- Becke (1996b) A. D. Becke, Canadian Journal of Chemistry 74, 995 (1996b).
- Becke (2002) A. D. Becke, The Journal of Chemical Physics 117, 6935 (2002).
- Maximoff et al. (2004) S. N. Maximoff, M. Ernzerhof, and G. E. Scuseria, The Journal of Chemical Physics 120, 2105 (2004).
- Burnus et al. (2005) T. Burnus, M. A. Marques, and E. K. Gross, Physical Review A 71, 010501 (2005).
- Pittalis et al. (2007) S. Pittalis, S. Kurth, S. Sharma, and E. K. U. Gross, The Journal of Chemical Physics 127, 124103 (2007).
- Pittalis et al. (2009) S. Pittalis, E. Räsänen, and E. K. U. Gross, Phys. Rev. A 80, 032515 (2009).
- Räsänen et al. (2009) E. Räsänen, S. Pittalis, C. R. Proetto, and E. K. U. Gross, Phys. Rev. B 79, 121305 (2009).
- Oliveira et al. (2010) M. J. T. Oliveira, E. Räsänen, S. Pittalis, and M. A. L. Marques, Journal of Chemical Theory and Computation 6, 3664 (2010).
- Zhu et al. (2016) W. Zhu, L. Zhang, and S. B. Trickey, The Journal of Chemical Physics 145, 224106 (2016), eprint http://aip.scitation.org/doi/pdf/10.1063/1.4971377, URL http://aip.scitation.org/doi/abs/10.1063/1.4971377.
- James W. Furness and Teale (2016) T. H. James W. Furness, Ulf Ekström and A. M. Teale, Molecular Physics 114, 1415 (2016), eprint https://doi.org/10.1080/00268976.2015.1133859, URL https://doi.org/10.1080/00268976.2015.1133859.
- Holzer et al. (2022) C. Holzer, Y. J. Franzke, and A. Pausch, The Journal of Chemical Physics 157, 204102 (2022), ISSN 0021-9606.
- not (c) SU(2) invariance of the exact xc-energy functional on time-reversal invariant states, implies Vignale and Rasolt (1988); Pittalis et al. (2017). The dependence of the exact xc-energy on the spin current via the covariant vorticity, Abedinpour et al. (2010) for the same class of states further implies . From the two relations, we readily obtain = 0.
- Aschebrock and Kümmel (2019) T. Aschebrock and S. Kümmel, Phys. Rev. Res. 1, 033082 (2019), URL https://link.aps.org/doi/10.1103/PhysRevResearch.1.033082.
- (61) Computational and implementation details are available at INSERT URL HERE.
- Desmarais et al. (2021a) J. K. Desmarais, S. Komorovsky, J.-P. Flament, and A. Erba, J. Chem. Phys. 154, 204110 (2021a).
- Desmarais et al. (2020b) J. K. Desmarais, J.-P. Flament, and A. Erba, Phys. Rev. B 101, 235142 (2020b).
- Desmarais et al. (2019a) J. K. Desmarais, J.-P. Flament, and A. Erba, J. Phys. Chem. Lett. 10, 3580 (2019a).
- Desmarais et al. (2019b) J. K. Desmarais, J.-P. Flament, and A. Erba, J. Chem. Phys. 151, 074107 (2019b).
- Metz et al. (2000) B. Metz, H. Stoll, and M. Dolg, J. Chem. Phys. 113, 2563 (2000).
- Peterson et al. (2003) K. A. Peterson, D. Figgen, E. Goll, H. Stoll, and M. Dolg, J. Chem. Phys. 119, 11113 (2003).
- Peterson et al. (2007) K. A. Peterson, D. Figgen, M. Dolg, and H. Stoll, J. Chem. Phys. 126 (2007).
- Doll et al. (2001) K. Doll, V. Saunders, and N. Harrison, Int. J. Q. Chem. 82, 1 (2001).
- Doll (2001) K. Doll, Comput. Phys. Commun. 137, 74 (2001).
- Doll et al. (2006) K. Doll, R. Dovesi, and R. Orlando, Theor. Chem. Acc. 115, 354 (2006).
- Civalleri et al. (2001) B. Civalleri, P. D’Arco, R. Orlando, V. Saunders, and R. Dovesi, Chem. Phys. Lett. 348, 131 (2001).
- Laun et al. (2018) J. Laun, D. Vilela Oliveira, and T. Bredow, J. Comp. Chem. 39, 1285 (2018).
- Heyd et al. (2005) J. Heyd, J. E. Peralta, G. E. Scuseria, and R. L. Martin, J. Chem. Phys. 123 (2005).
- Lebedev (1976) V. I. Lebedev, USSR Comput. Math. Math. Phys. 16, 10 (1976).
- Lebedev (1977) V. I. Lebedev, Sib. Math. J. 18, 99 (1977).
- Towler et al. (1996) M. D. Towler, A. Zupan, and M. Causà, Comput. Phys. Commun. 98, 181 (1996).
- Lehtola et al. (2020) S. Lehtola, F. Blockhuys, and C. Van Alsenoy, Molecules 25 (2020).
- Furness et al. (2022) J. W. Furness, A. D. Kaplan, J. Ning, J. P. Perdew, and J. Sun, J. Chem. Phys. 156 (2022).
- (80) RPAx calculations were performed with NWCHEM version 7.0, see https://nwchemgit.github.io/Excited-State-Calculations.html.
- Aprà et al. (2020) E. Aprà, E. J. Bylaska, W. A. de Jong, N. Govind, K. Kowalski, T. P. Straatsma, M. Valiev, H. J. J. van Dam, Y. Alexeev, J. Anchell, et al., J. Chem. Phys. 152, 184102 (2020).
- Desmarais et al. (2021b) J. K. Desmarais, A. Erba, J.-P. Flament, and B. Kirtman, J. Chem. Theor. Comput. 17, 4697 (2021b).
- Desmarais et al. (2021c) J. K. Desmarais, A. Erba, J.-P. Flament, and B. Kirtman, J. Chem. Theor. Comput. 17, 4712 (2021c).
- Desmarais et al. (2023b) J. K. Desmarais, A. Boccuni, J.-P. Flament, B. Kirtman, and A. Erba, J. Chem. Theor. Comput. 19, 1853 (2023b).
- Ullrich (2011) C. A. Ullrich, Time-dependent density-functional theory: concepts and applications (OUP Oxford, 2011).
- Desmarais et al. (2023c) J. K. Desmarais, A. Erba, and J.-P. Flament, Phys. Rev. B 108, 134108 (2023c).
- Gulans and Draxl (2022) A. Gulans and C. Draxl, arXiv preprint arXiv:2204.02751 (2022).
- Choi et al. (2017) B. K. Choi, M. Kim, K.-H. Jung, J. Kim, K.-S. Yu, and Y. J. Chang, Nanoscale Res. Lett. 12, 1 (2017).
- Island et al. (2016) J. O. Island, A. Kuc, E. H. Diependaal, R. Bratschitsch, H. S. Van Der Zant, T. Heine, and A. Castellanos-Gomez, Nanoscale 8, 2589 (2016).
- Zelewski and Kudrawiec (2017) S. J. Zelewski and R. Kudrawiec, Sci. Reports 7, 15365 (2017).
- Conan et al. (1979) A. Conan, D. Delaunay, A. Bonnet, A. Moustafa, and M. Spiesser, physica status solidi (b) 94, 279 (1979).
- Oliva et al. (2020) R. Oliva, T. Woźniak, F. Dybala, J. Kopaczek, P. Scharoch, and R. Kudrawiec, Mat. Res. Lett. 8, 75 (2020).
- Zhang et al. (2014) Y. Zhang, T.-R. Chang, B. Zhou, Y.-T. Cui, H. Yan, Z. Liu, F. Schmitt, J. Lee, R. Moore, Y. Chen, et al., Nature nano. 9, 111 (2014).
- Shim et al. (2014) G. W. Shim, K. Yoo, S.-B. Seo, J. Shin, D. Y. Jung, I.-S. Kang, C. W. Ahn, B. J. Cho, and S.-Y. Choi, ACS nano 8, 6655 (2014).
- Ross et al. (2013) J. S. Ross, S. Wu, H. Yu, N. J. Ghimire, A. M. Jones, G. Aivazian, J. Yan, D. G. Mandrus, D. Xiao, W. Yao, et al., Nature Commun. 4, 1474 (2013).
- Beal et al. (1972) A. Beal, J. Knights, and W. Liang, J. Phys. C Sol. State Phys. 5, 3540 (1972).
- Ruppert et al. (2014) C. Ruppert, B. Aslan, and T. F. Heinz, Nano Lett. 14, 6231 (2014).
- Skone et al. (2014) J. H. Skone, M. Govoni, and G. Galli, Phys. Rev. B 89, 195112 (2014).
- Lv et al. (2015) B. Lv, H. Weng, B. Fu, X. P. Wang, H. Miao, J. Ma, P. Richard, X. Huang, L. Zhao, G. Chen, et al., Phys. Rev. X 5, 031013 (2015).
- Huang et al. (2015) S.-M. Huang, S.-Y. Xu, I. Belopolski, C.-C. Lee, G. Chang, B. Wang, N. Alidoust, G. Bian, M. Neupane, C. Zhang, et al., Nat. Comms. 6, 7373 (2015).
- Jana et al. (2018) S. Jana, A. Patra, and P. Samal, The Journal of Chemical Physics 149, 044120 (2018), ISSN 0021-9606, eprint https://pubs.aip.org/aip/jcp/article-pdf/doi/10.1063/1.5040786/14927796/044120_1_online.pdf, URL https://doi.org/10.1063/1.5040786.
- Ekholm et al. (2018) M. Ekholm, D. Gambino, H. J. M. Jönsson, F. Tasnádi, B. Alling, and I. A. Abrikosov, Phys. Rev. B 98, 094413 (2018), URL https://link.aps.org/doi/10.1103/PhysRevB.98.094413.
- Fu and Singh (2018) Y. Fu and D. J. Singh, Phys. Rev. Lett. 121, 207201 (2018), URL https://link.aps.org/doi/10.1103/PhysRevLett.121.207201.
- Fu and Singh (2019) Y. Fu and D. J. Singh, Phys. Rev. B 100, 045126 (2019), URL https://link.aps.org/doi/10.1103/PhysRevB.100.045126.
- Mejía-Rodríguez and Trickey (2019) D. Mejía-Rodríguez and S. B. Trickey, Phys. Rev. B 100, 041113 (2019), URL https://link.aps.org/doi/10.1103/PhysRevB.100.041113.
- Romero and Verstraete (2018) A. Romero and M. Verstraete, Eur. Phys. J. B 91, 193 (2018), URL https://link.springer.com/article/10.1140/epjb/e2018-90275-5#citeas.
- Buchelnikov et al. (2019) V. D. Buchelnikov, V. V. Sokolovskiy, O. N. Miroshkina, M. A. Zagrebin, J. Nokelainen, A. Pulkkinen, B. Barbiellini, and E. Lähderanta, Phys. Rev. B 99, 014426 (2019), URL https://link.aps.org/doi/10.1103/PhysRevB.99.014426.
- Tran et al. (2020) F. Tran, G. Baudesson, J. Carrete, G. K. H. Madsen, P. Blaha, K. Schwarz, and D. J. Singh, Phys. Rev. B 102, 024407 (2020), URL https://link.aps.org/doi/10.1103/PhysRevB.102.024407.